A sample of nitrogen gas occupies 1.55 L at 27.0°C and 1.00 atm. What will ths volume be at -100.0°C with the same pressure?
1
Expert's answer
2020-05-07T14:15:57-0400

Charles Law: V1/T1=V2/T2;

T1=273+27=300K; T2=273-100=173K

V2=V1*T2/T1;

V2=1.55*173/300=0.894 L

Answer: 0.894 L